Please help
The domain of the function f.x --> x^3-5 is D={0,1,2,3,4}. Find the range of f

Answer by stanbon(75887)

The domain of the function f.x --> x^3-5 is D={0,1,2,3,4}. Find the range of f
----------------
f(0) = 0^3-5 = -5
f(1) = 1^3-5 = -4
f(2) = 2^3-5 = 3
f(3) = 3^3-5 = 22
f(4) = 4^3-5 = 59
---------
Range:: {-5,-4,3,22,59}
========================
Cheers,
Stan H.
---------------------